OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a15 = -53, a16 = -5 a. an = -725 - 48(n - 1) b. an = -725 + 48(n + 1) c. an = -725 + 48(n - 1) d. an = -725 - 48(n + 1)

hartnn (hartnn):

find the common difference(d) first, its the difference between the consecutive terms \(\large d = a_{16}-1_{15} = ... ? \)

hartnn (hartnn):

i meant \(\large d = a_{16}-a_{15} = ... ?\)

OpenStudy (anonymous):

48

hartnn (hartnn):

thats correct, d = 48 now we just need first term a1 use the general n'th term formula \(a_n = a_1 + (n-1)d \\ a_{16} = -5 = a_1+(16-1)\times 48\) can you find a1 from there ?

OpenStudy (anonymous):

-725

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

so is the answer an = -725 + 48(n - 1)

hartnn (hartnn):

yes it is! :)

OpenStudy (anonymous):

thxx

hartnn (hartnn):

wlcmxx